find the values ot the trigonometric functions of the angle ( in standard position) whose terminal side passes through the point (24,7).
Answer by stanbon(75887)

find the values ot the trigonometric functions of the angle ( in standard position) whose terminal side passes through the point (24,7).
-------
y = 7
x = 24
-----
So, r = sqrt[7^2+24^2] = sqrt(625) = 25
------
sin = y/r = 7/25
cos = x/4 = 24/25
tan = y/x = 7/24
----
csc, sec, cot are the inverse of sin,cos,tan
------------------------------------------------
Cheers,
Stan H.
----------------